# Some insights into depth estimators for location and scatter in the multivariate setting

### v1

@@ Abstract
The concept of statistical depth has received considerable attention as a way to extend the notions of the median and quantiles to other statistical models. These procedures aim to formalize the idea of identifying deeply embedded fits to a model that are less influenced by contamination. Since contamination introduces bias in estimators, it is well known in the location model that the median minimizes the worst-case performance, in terms of maximum bias, among all equivariant estimators. In the multivariate case, Tukey’s median was a groundbreaking concept for location estimation, and its counterpart for scatter matrices has recently attracted considerable interest. The breakdown point and the maximum asymptotic bias are key concepts used to summarize an estimator’s behavior under contamination. For the location and scale model, we consider two closely related depth formulations, whose deepest estimators display significantly different behavior in terms of breakdown point. In the multivariate setting, we analyze recently introduced concentration inequalities that provide a unified framework for studying both the statistical convergence rate and robustness of Tukey’s median and depth-based scatter matrices. We observe that slight variations in these inequalities allow us to visualize the maximum bias behavior of the deepest estimators. Since the maximum bias for depth-based scatter matrices had not previously been derived, we explicitly calculate both the breakdown point and the maximum bias curve for the deepest scatter matrices.
@@ 1 Introduction
The concept of statistical depth has received considerable attention as a way to extend the notions of the median and quantiles to more general statistical models. Originally introduced by Tukey (1975), depth was defined as the minimum proportion of data points lying on either side of a point, and this idea was later extended to the bivariate case. Donoho and Gasko (1992) formalized the general notion of the depth of a point $\mathbf{z}$ in a $p$-dimensional space with respect to a probability measure $P$.
@@
The seminal paper by Huber (1964) highlighted the median as an estimator with optimal worst-case behavior under contamination (i.e., its maximum bias), making it the most robust among the class of equivariant location estimators. A key feature of the median is that it is flanked on each side by half of the data mass, helping to shield it from the effects of outliers. This robustness—being well-surrounded by data—is echoed in Tukey’s concept of the multivariate median. In this case, a point in Euclidean space is considered a median if it maximizes the minimum mass contained in any closed half-space whose boundary includes the point.
@@
This pursuit of a point that is ”deeply inserted” into the data, to enhance robustness in univariate and multivariate location models, has been extended to other statistical frameworks. In the regression setting, Rousseeuw and Hubert (1999) defined regression depth, which measures how deeply a linear fit is embedded within the data. This is evaluated by the smallest amount of data mass in the two opposing wedges formed by the fit plane and vertical planes orthogonal to the explanatory subspace. Mizera and Müller (2004) further explored depth concepts in the context of location-scale models.
@@
Chen and Tyler (2002) investigated various properties of Tukey’s median, including its influence function and maximum contamination bias. Chen et al. (2018) extended the idea of depth to covariance matrix estimation by introducing the concept of matrix depth. Their estimator achieves the optimal rate under Huber’s $\varepsilon$-contamination model for estimating covariance/scatter matrices with various structural assumptions, such as bandedness and sparsity. Paindaveine and Van Bever (2018) also developed halfspace depth concepts for scatter, concentration, and shape matrices. While their concept of scatter depth is similar to that in Chen et al. (2018) and Zhang (2002), rather than focusing on the deepest scatter matrix, they studied the properties of the depth function and its associated depth regions.
Nagy et al. (2019) discussed the relationships between halfspace depth and problems in affine and convex geometry, offering an extensive overview of various depth notions.
@@
Section 2 presents two proposals for joint location and scale depth estimation in the univariate location-scale model, comparing their differing breakdown behaviors despite conceptual similarity. Section 3 discusses the importance of maximum asymptotic bias in understanding how estimator behavior evolves as sample size $n$ becomes large. By examining the concentration inequalities in Chen et al. (2018), we show that slight variations in these inequalities allow us to visualize the maximum bias behavior of the deepest estimators. In particular, Section 4 derives the maximum bias function for the deepest scatter matrix estimator introduced by Chen et al. (2018) in the case of known location, showing that this estimator shares Tukey’s median’s asymptotic breakdown point of $1/3$. All proofs are deferred to the Appendix.
2 The location-scale model, depth and breakdown point
Let us take the multivariate location and scatter model,.
$X\mathbf{=\mu}_{0}\mathbf{+V}_{0}\mathbf{u}$ (1)
@@ 2 The location-scale model, depth and breakdown point
with $\mathbf{\mu}_{0}\in\mathbb{R}^{p},$ $\mathbf{V}_{0}\in\mathbb{R}^{p\times p}$ an inversible matrix, and $\mathbf{u}$ an inobservable random vector with centrosymmetric distribution $P_{\mathbf{0}}$ around $\mathbf{0}.$ $\mathbf{\mu}_{0}$ and $\mathbf{V}_{0}\mathbf{V}_{0}^{t}$ (except for a constant) are identifiable. Tukey’s depth of a vector $\mathbf{\theta}\in\mathbb{R}^{p}$ is given by

### v2

@@ Abstract
The concept of statistical depth extends the notions of the median and quantiles to other statistical models. These procedures aim to formalize the idea of identifying deeply embedded fits to a model that are less influenced by contamination. In the multivariate case, Tukey’s median was a groundbreaking concept for multivariate location estimation, and its counterpart for scatter matrices has recently attracted considerable interest. The breakdown point and the maximum asymptotic bias are key concepts used to summarize an estimator’s behavior under contamination. In the multivariate and regression setting we analyze recently introduced concentration inequalities that provide a unified framework for studying both the statistical convergence rate and robustness of Tukey’s median, depth-based scatter matrices and depth-based multivariate regression estimators. We observe that slight variations in these inequalities allow us to visualize the maximum bias behavior of the deepest estimators. We explicitly obtain the maximum bias curve and breakdown point of the deepest scatter matrices. For the location and scale model, we consider two closely related depth formulations, whose deepest estimators display significantly different behavior in terms of breakdown point. A numerical study is performed to compare the finite sample bias performance of several robust estimators in the multivariate setting.
@@ 1 Introduction
The concept of statistical depth has received considerable attention as a way to extend the notions of the median and quantiles to more general statistical models. Originally introduced by Tukey (1975), depth was defined as the minimum proportion of data points lying on either side of a point, and this idea was later extended to the bivariate case. Donoho and Gasko (1992) formalized the general notion of the halfspace depth of a point $\mathbf{z}$ in a $p$-dimensional space with respect to a probability measure $P$ in which the search is for a location or center of the probability measure $P$.
@@
The seminal paper by Huber (1964) highlighted the median as an estimator with optimal worst-case behavior under contamination (i.e., its maximum bias), making it the most robust among the class of equivariant location estimators. A key feature of the median is that it is flanked on each side by half of the data mass, helping to shield it from the effects of outliers. This robustness—being well-surrounded by data—is echoed in Tukey’s concept of the multivariate median. In this case, a point in Euclidean space is considered a median if it maximizes the minimum mass contained in any closed half-space whose boundary includes the point. In this sense, the search for a point or location center that fits deeply within the data cloud is the idea that pervades some other statistical models.
@@
This pursuit of a point that is ”deeply inserted” into the data, to enhance robustness in univariate and multivariate location models, has been extended to other statistical frameworks. In the regression setting, Rousseeuw and Hubert (1999) defined regression depth, which measures how deeply a linear fit is embedded within the data. This is evaluated by the smallest amount of data mass in the two opposing wedges formed by the fit plane and vertical planes orthogonal to the explanatory subspace. Mizera and Müller (2004) further explored depth concepts in the context of location-scale models. In fact, the univariate regression depth treated by Rousseeuw and Hubert (1999) is a special case of the multivariate regression depth defined by Mizera (2002). A very similar but earlier definition was proposed in Bern and Eppstein (2000).
@@
Chen, Gao and Ren (2018a) extended the idea of depth to covariance matrix estimation by introducing the concept of matrix depth. Their estimator achieves the optimal rate under Huber’s $\varepsilon$-contamination model for estimating covariance/scatter matrices with various structural assumptions, such as bandedness and sparsity. Paindaveine and Van Bever (2018) also developed halfspace depth concepts for scatter, concentration, and shape matrices. While their concept of scatter depth is similar to that of Chen, Gao and Ren (2018a) and Zhang (2002), rather than focusing on the deepest scatter matrix, they studied the properties of the depth function and its associated depth regions. Gao (2020) considered estimators that are maximizers of multivariate regression depth functions, by studying its minimax rates in the settings of $\varepsilon$-contamination models for various regression problems including nonparametric regression, sparse linear regression, reduced rank regression, etc. Nagy, Schütt and Werner (2019) discussed the relationships between halfspace depth and problems in affine and convex geometry, offering an extensive overview of various depth notions.
The concept of asymptotic maximum bias for estimators in a neighborhood has been treated in the robust literature since the pioneering paper by Huber (1964). The maximum asymptotic bias function depicts more accurately the behavior of the estimator since it captures the global behavior by quantifying how much the estimator deviates in the whole contamination neighborhood under different levels of contamination from the parameters at the central model. Among several contributions in the context of maximum bias we can mention: He and Simpson (1993) obtained a lower bound on the contamination bias of an estimator that holds for a wide class of parametric families and a second lower bound which applies to locally linear estimates yielding that such estimates cannot be bias minimax among all Fisher-consistent estimates in higher dimensions; Martin and Zamar (1989) derived asymptotically minimax bias robust M-estimates of scale for positive random variables in $\varepsilon$-contamination neighborhoods; Martin and Zamar (1993a) dealt with minimax location M-estimators under an efficiency constraint; Martin and Zamar (1993b) established expressions for the maximum asymptotic bias of M-estimates of scale over the $\varepsilon$-contamination neighborhood and they construct asymptotically minimax bias robust estimates of scale; Ghement, Ruiz and Zamar (2008) considered robust scale estimators in the nonparametric regression setting and they derived their maximum bias curves; Martin, Yohai and Zamar (1989) found minimax bias estimators for two different classes of regression estimates: M-estimates with general scale and GM-estimates in the context of known intercept and elliptical regressors; Maronna and Yohai (1993) delved into a projection estimator for regression whose maximum bias is twice the lower bound attainable for maximum bias in regression; Berrendero and Zamar (2001) calculated the maximum bias for several classes of estimators in a general setup of unknown intercept and nonelliptical regressors; Svarc, Yohai and Zamar (2002) found a family of M-estimates of regression such that they minimize the asymptotic variance at the central model subject to a bound on the maximum bias over contamination neighborhoods; Yohai and Maronna (1990) calculated the maximum bias for the minimum volume ellipsoid and Tyler scatter estimator under point mass contaminations; Maronna, Stahel and Yohai (1992) came up with a projection estimator for multivariate scatter with remarkable bias performance; Maronna and Yohai (1995) studied the bias behavior of the Stahel-Donoho estimator for scatter among a bundle of multivariate dispersion estimators, Adrover (1998) computed the asymptotic maximum bias for M-estimators for multivariate dispersion showing that Tyler scatter M-estimator minimizes the maximum bias in the $\varepsilon$-neighborhood; Chen and Tyler (2002) investigated various properties of Tukey’s median, including its influence function and maximum contamination bias; Adrover, Maronna and Yohai (2002) derived the maximum asymptotic bias for the regression depth estimator and compared its performance with that of Maronna and Yohai (1993); Zuo, Cui and Young (2004) computed the maximum bias and influence functions for projection based estimators for multivariate location; Zamar (1992) calculated the maximum bias of M-estimators for orthogonal regression over the $\varepsilon$-contamination neighborhood and obtained the M-estimator minimizing the maximum bias in such a class.
@@
Section 2 reviews the concept of halfspace depth in various settings: multivariate location and scatter, regression and multivariate regression models. Section 3 discusses the importance of maximum asymptotic bias in understanding how the estimator behavior evolves as sample size $n$ becomes sufficiently large. By examining the concentration inequalities in Chen, Gao and Ren (2018a), we show that slight variations in the derivation of these inequalities allow us to visualize the maximum bias behavior of the deepest estimators. In particular, Section 4 derives the maximum bias function for the deepest scatter matrix estimator introduced by Chen, Gao and Ren (2018a) in the case of known location, showing that this estimator shares Tukey’s median’s asymptotic breakdown point of $1/3$. Section 5 deals with two proposals for joint location and scale depth estimation in the univariate location-scale model, comparing their differing breakdown behaviors despite conceptual similarity. In Section 6 a simulation analysis is performed including several robust proposals to compare their performance in the presence of contamination, when the location estimator is included. All proofs are deferred to the Appendix.
2 Depth in several statistical models: a review
Let us consider the multivariate location and scatter model,
$\mathbf{X=\mu}_{0}\mathbf{+}V_{0}\mathbf{u}$ (1)
@@ 2 Depth in several statistical models: a review
with $\mathbf{\mu}_{0}\in\mathbb{R}^{p},$ $V_{0}\in\mathbb{R}^{p\times p}$ an invertible matrix, and $\mathbf{u\sim}P_{0}$ an unobservable random vector. $\mathbf{\mu}_{0}$ and $V_{0}V_{0}^{t}$ (except for a constant) are assumed to be identifiable (for instance this holds if $P_{0}$ is a centrosymmetric distribution around $\mathbf{0)}$. Tukey’s depth of a vector $\mathbf{\theta}\in\mathbb{R}^{p}$ is given by
$D_{T}\left(\mathbf{\theta},P\right)=\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}P\left(\mathbf{u}^{t}\mathbf{X}\leq\mathbf{u}^{t}\mathbf{\theta}\right),$ (2)